# Brenda Prallon22footnotemark: 2    Chen Qiu22footnotemark: 2    Yiwei Sun22footnotemark: 2

### v1

@@ B.1 Gurobi’s Output
In this section, we discuss some parameters in the Gurobi optimizer that can be tuned to improve its performance. [^21] For more specifics, see https://www.gurobi.com/documentation/current/refman/mip_models.html. First, MIPFocus specifies, on a high level, whether the solution should prioritize speed or optimality. The default value of this parameter is 00, which achieves a balance between searching for new solutions and proving optimality of the current solution. We set this parameter to $2$, prioritizing finding good-quality and optimal solutions. A second set of important parameters affect how the solver terminates. There are two termination choices for MIP models: 1) a restriction on the runtime of the algorithm, such as using TimeLimit to limit the wall-clock runtime, and 2) controlling the optimality gap, by setting a parameter MIPGap that stops the algorithm when the relative gap between the best-known solution and the best known bound on the solution objective is less than the specified value. In this application, we let the runtime to be the default value (infinity) and set the tolerance level to $10^{-6}$. In Figure 9, the Gurobi solver outputs not only the solution and the optimal value of the problem but also the output gap. In this example, we can see that the gap between bounds is 00, demonstrating that the $k$-median problem has been solved to provable optimality. An interesting feature of using the integer programming formulation of the $k$-median problem is that even when we are not able to solve the problem to provable optimality, the optimality gap provides a suboptimality guarantee for the obtained solution. Finally, the output reports the algorithm’s runtime, which was only $0.05$ seconds for this example. There are more parameters one can tune to control the root relaxation, the aggressiveness of the cutting plane strategies and the level of presolve, tolerance parameters for primal feasibility, the integer feasibility, and more. For the applications in this paper, we let the rest of the parameters be the default.
@@
Consider the following specialization of our setting: $d=1$ and therefore $X\in\mathbb{R}$, $\mathcal{S}_{E}=\{1,4\}$, $\mathcal{S}_{P}=\{2,3\}$, and for all sites we just have $X_{s}=s$. (Intuitively, we consider 4 sites lined up equidistantly on a straight line, where the middle two sites are the policy sites.) Suppose furthermore that sampling distributions of signals are degenerate; formally, $\hat{\tau}_{s}=\tau_{s}:=\tau(X_{s})$ for all $s$.
@@
    If sampling has to be purposive (nonrandomized), the lowest regret achievable in combination with any treatment rule equals $43C/$ 4.
@@
Next, suppose the policy maker uniformly randomizes over $\mathscr{S}\subset\{1,4\}$ for experimentation and then implements the new policy with probability $a_{2}=a_{3}=[(C+\hat{\tau}_{\mathscr{S}})/2C]_{0}^{1}$; here, the notation $[X]_{0}^{1}:=\min\{\max\{X,0\},1\}$ indicates clamping of $X$ to $[0,1]$. We will show that the worst-case expected regret under this scheme is $C/2$, which therefore is the problem’s MMR value and is attained by this rule.
@@
We will solve this by considering subcases. Suppose first that $\tau_{2}$ and $\tau_{3}$ have different signs. Since both objective and constraints are invariant under multiplying $(\tau_{1},\tau_{2},\tau_{3},\tau_{4})$ by $-1$, suppose without further loss of generality that $\tau_{2}\geq 0\geq\tau_{3}$. The optimization problem now simplifies to
@@
Clamping of expressions at $1$ cannot bind because $\tau_{2}$ and $\tau_{3}$ are positive. If clamping at 0 binds for both fractions, then the objective equals 00. Suppose clamping at 00 binds for one expression, say (without further loss of generality) because $\tau_{2}>2C$, then we have simplification to
@@
Keeping in mind that $\tau_{2}>2C$ and therefore also $\tau_{3}>C$ in this subcase, evaluation of derivatives shows that this expression decreases in $\tau_{3}$; hence, $\tau_{3}=\tau_{2}-C$. Substituting this in, one can further verify the expression to be decreasing in $\tau_{2}$; therefore, the maximal value in this subcase is attained at a boundary point also covered by the next case (and, though not essential for the argument, this value can be verified to be $83C/$ 8).
@@
Regarding claim 2, by the decision problem’s symmetry, it is without further loss of generality to assume that site $1$ is being sampled. MMR is then at least $43C/$ 4 because this value is achieved if the true parameter values $(\tau_{1},\tau_{2},\tau_{3},\tau_{4})$ are equally likely to be $(0,C,2C,3C)$ or $-(0,C,2C,3C)$.
We next show that this value is attained by uniformly assigning treatment with probability $a_{2}=a_{3}=[(3C-2\hat{\tau}_{1})/6C]_{0}^{1}$. Indeed, worst case regret of this decision rule equals
@@
If $\tau_{2}$ and $\tau_{3}$ have different signs, we can bound this value by $C/2$ just as before. Suppose they have the same sign, which we take to be positive without further loss of generality, then the problem simplifies to
@@
Here, the first equality concentrates out $\tau_{2}=\tau_{1}+C$ and $\tau_{3}=\tau_{2}+C$; the second equality uses that clamping cannot bind (clamping at 00 would set the expression to 00, clamping at $1$ would imply that $\tau_{2}<0$), after which the problem is straightforwardly solved by $\tau_{1}=0$. ∎
@@ B.3 Additional Analysis of the Survey Experiment
As mentioned in Section 5, we can obtain experimental estimates of all fifteen countries using the original experiments conducted in Naumann et al. (2018). Generally, experimental estimates of all policy sites are unknown and unattainable in most real-world applications; otherwise, there is no need to solve the site selection problem. However, as an illustrative example, we will leverage the information in these experiments to quantify the magnitude of the Lipschitz constant $C$ needed to explain the treatment heterogeneity in the data and the constant $C(\mathscr{S})$ in the assumption of Lemma 1 that gives the result of this paper.
The outcome of interest from the survey is a categorical variable indicating survey respondents’ attitudes towards immigrants: $1$ for Allow None; $2$ for Allow A Few; $3$ for Allow Some; $4$ for Allow Many. For a more straightforward interpretation, we redefine the outcome variable to be binary: we let the outcome of the survey respondent be $1$ if she answers $3$ or $4$, indicating “support”; otherwise, we let her outcome be 00, indicating “oppose.” The treatment is also a binary variable, which equals $1$ $(0)$ if the survey is about high-skilled (low-skilled) immigrants. We use a simple difference-in-means estimator to estimate the treatment effect of each country. The table below shows the point estimates and their standard errors. The point estimates speak to the difference between the percentage of people who support high-skilled immigrants and the percentage of people who support low-skilled immigrants. For example, 25.8% more survey respondents in Austria are more supportive of high-skilled immigrants, compared to low-skilled immigrants.
@@
Notes: This table presents the experimental estimates of the treatment effect of the policy in each country. The original outcome from the survey is a categorical variable indicating survey respondents’ attitudes towards immigrants: $1$ for Allow None; $2$ for Allow A Few; $3$ for Allow Some; $4$ for Allow Many. We redefine the outcome variable to be binary: we let the outcome of the survey respondent be $1$ if she answers $3$ or $4$, indicating “support”; otherwise, we let her outcome be 00, indicating “oppose.” The treatment is also a binary variable, which equals $1$ $(0)$ if the survey is about high-skilled (low-skilled) immigrants. We use a simple difference-in-means estimator to estimate the treatment effect of each country and their standard error.
@@
$\max\left\{\frac{|\hat{\tau}(X_{i})-\hat{\tau}(X_{j})|}{\|X_{i}-X_{j}\|}\right\},\quad\forall i,j\in\mathcal{S}_{E}\cup\mathcal{S}_{P},$ (44)
@@
Notes: The dots in Panel 10(a) represent all possible pairs among the $15$ countries. The slope of each point connecting to the origin represents the value of the Lipschitz constant needed to explain the estimated treatment effects and site-level covariates observed for that pair of countries. The red dash line is the smallest $C$ that is needed to explain all the data, computed using equation 44. The slope of the blue dashed line represents $C(\mathscr{S})$, computed using 45. Panel 10(b) presents a histogram of $C(\mathscr{S})$ for all possible choices of experimental sites $\mathscr{S}$. The red dashed line is the smallest Lipschitz constant needed to explain the data, corresponding to the same red dashed line in panel 10(a).
@@ B.4 A Simple Example that Motivates Assumption 4
where $X_{i,s}\in\mathbb{R}$ is the observed unit-level covariate for individual $i$ in site $s$, $Z_{i,s}\in\mathbb{R}$ is the unobserved unit-level covariate, $\varepsilon_{i,s}\sim\mathcal{N}(0,\sigma_{\varepsilon,s}^{2})$ is an error term with $\sigma_{\varepsilon,s}^{2}>0$ and is also independent of $\left(X_{i,s},Z_{i,s}\right)$, and $\beta,\gamma\in\mathbb{R}$ are the same across different sites. For simplicity, suppose that $X_{i,s}$ and $Z_{i,s}$ are jointly normal: $\left(X_{i,s},Z_{i,s}\right)^{\top}\sim\mathcal{N}(\mu_{s},\varSigma_{s})$, where $\mu_{s}\in\mathbb{R}^{2}$ and $\varSigma_{s}$ is a $2\times 2$ positive-definite covariance matrix.
Let $\bar{Y}_{s}:=\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}Y_{i,s}$ be the sample average of the observed outcome at site $s$, and $\bar{X}_{s}:=\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}X_{i,s},\bar{Z}_{s}:=\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}Z_{i,s}$ be the observed and unobserved site-level aggregate covariates, respectively. Under the above assumptions and conditional on observed site-level aggregate covariate $\bar{X}_{s}$, we have
@@
where $\mathbb{E}[\bar{Z}_{s}|\bar{X}_{s}]$ is the expectation of $\bar{Z}_{s}$ conditional $\bar{X}_{s}$, and $\sigma_{s}^{2}>0$. Then, Assumption 4 models a case for which the policy effect of interest is
@@
Furthermore, (46) implies that, conditional on $\bar{X}_{s}$, we have an unbiased and normal estimator for $\tau_{s}$. We may also calculate that $\mathbb{E}[\bar{Z}_{s}|\bar{X}_{s}]=\alpha_{s}$ for some $\alpha_{s}$ that depends on each site $s$. Then, for $s,s^{\prime}\in\mathcal{S}$, $s\neq s^{\prime}$, we have:

### v2

@@ B.1 Gurobi’s Output
In this section, we discuss some parameters in the Gurobi optimizer that can be tuned to improve its performance. [^22] For more specifics, see https://www.gurobi.com/documentation/current/refman/mip_models.html. First, MIPFocus specifies, on a high level, whether the solution should prioritize speed or optimality. The default value of this parameter is 0, which achieves a balance between searching for new solutions and proving optimality of the current solution. We set this parameter to $2$, prioritizing finding good-quality and optimal solutions. A second set of important parameters affect how the solver terminates. There are two termination choices for MIP models: 1) a restriction on the runtime of the algorithm, such as using TimeLimit to limit the wall-clock runtime, and 2) controlling the optimality gap, by setting a parameter MIPGap that stops the algorithm when the relative gap between the best-known solution and the best known bound on the solution objective is less than the specified value. In this application, we let the runtime to be the default value (infinity) and set the tolerance level to $10^{-6}$. In Figure 9, the Gurobi solver outputs not only the solution and the optimal value of the problem but also the output gap. In this example, we can see that the gap between bounds is 0, demonstrating that the $k$-median problem has been solved to provable optimality. An interesting feature of using the integer programming formulation of the $k$-median problem is that even when we are not able to solve the problem to provable optimality, the optimality gap provides a suboptimality guarantee for the obtained solution. Finally, the output reports the algorithm’s runtime, which was only $0.05$ seconds for this example. There are more parameters one can tune to control the root relaxation, the aggressiveness of the cutting plane strategies and the level of presolve, tolerance parameters for primal feasibility, the integer feasibility, and more. For the applications in this paper, we let the rest of the parameters be the default.
@@
Consider the following specialization of our setting: $d=1$ and therefore $X\in\mathbb{R}$, $\mathcal{S}_{E}=\{1,4\}$, $\mathcal{S}_{P}=\{2,3\}$, and for all sites we just have $X_{s}=s$. (Intuitively, we consider 4 sites lined up equidistantly on a straight line, where the middle two sites are the policy sites.) Suppose furthermore that sampling distributions of signals are degenerate; formally, $\widehat{\tau}_{s}=\tau_{s}:=\tau(X_{s})$ for all $s$.
@@
    If sampling has to be purposive (nonrandomized), the lowest regret achievable in combination with any treatment rule equals $3C/4$.
@@
Next, suppose the policymaker uniformly randomizes over $\mathscr{S}\subset\{1,4\}$ for experimentation and then implements the new policy with probability $a_{2}=a_{3}=[(C+\widehat{\tau}_{\mathscr{S}})/2C]_{0}^{1}$; here, the notation $X]_{0}^{1}:=\min\{\max\{X,0\},1\}$ indicates clamping of $X$ to $[0,1]$, and we denote $\tau_{s}^{+}:=\max\{\tau_{s},0\}$, $\tau_{s}^{-}:=\min\{\tau_{s},0\}$. We will show that worst-case regret under this scheme is $C/2$, which therefore is the problem’s MMR value and is attained by this rule.
@@
We will solve this by considering subcases. Suppose first that $\tau_{2}$ and $\tau_{3}$ have different signs. Since both objective and constraints are invariant under multiplying $(\tau_{1},\tau_{2},\tau_{3},\tau_{4})$ by 1-1, suppose without further loss of generality that $\tau_{2}\geq 0\geq\tau_{3}$. The optimization problem now simplifies to
@@
Clamping of expressions at $1$ cannot bind because $\tau_{2}$ and $\tau_{3}$ are positive. If clamping at 0 binds for both fractions, then the objective equals 0. Suppose clamping at 0 binds for one expression, say (without further loss of generality) because $\tau_{2}>2C$, then we have simplification to
@@
Keeping in mind that $\tau_{2}>2C$ and therefore also $\tau_{3}>C$ in this subcase, evaluation of derivatives shows that this expression decreases in $\tau_{3}$; hence, $\tau_{3}=\tau_{2}-C$. Substituting this in, one can further verify the expression to be decreasing in $\tau_{2}$; therefore, the maximal value in this subcase is attained at a boundary point also covered by the next case (and, though not essential for the argument, this value can be verified to be $3C/8$).
@@
Regarding claim 2, by the decision problem’s symmetry, it is without further loss of generality to assume that site $1$ is being sampled. MMR is then at least $3C/4$ because this value is achieved if the true parameter values $(\tau_{1},\tau_{2},\tau_{3},\tau_{4})$ are equally likely to be $(0,C,2C,3C)$ or (0,C,2C,3C)-(0,C,2C,3C).
We next show that this value is attained by uniformly assigning treatment with probability $a_{2}=a_{3}=[(3C-2\widehat{\tau}_{1})/6C]_{0}^{1}$. Indeed, worst case regret of this decision rule equals
@@
If $\tau_{2}$ and $\tau_{3}$ have different signs, we can bound this value by $C/2$ just as before. If they have the same sign, taken to be positive without further loss of generality, then the problem simplifies to
@@
Here, the first equality concentrates out $\tau_{2}=\tau_{1}+C$ and $\tau_{3}=\tau_{2}+C$; the second equality uses that clamping cannot bind (clamping at 0 would set the expression to 0, clamping at $1$ would imply that $\tau_{2}<0$), after which the problem is straightforwardly solved by $\tau_{1}=0$. ∎
@@ B.3 Additional Analysis of the Survey Experiment
As mentioned in Section 5.2, we can obtain experimental estimates of all fifteen countries using the original experiments conducted in Naumann et al. (2018). Generally, experimental estimates of all policy sites are unknown and unattainable in most real-world applications; otherwise, there is no need to solve the site selection problem. However, as an illustrative example, we will leverage the information in these experiments to quantify the magnitude of the Lipschitz constant $C$ needed to explain the treatment heterogeneity in the data and the constant $C(\mathscr{S})$ in the assumption of Lemma 1 that gives the result of this paper.
The outcome of interest from the survey is a categorical variable indicating survey respondents’ attitudes towards immigrants: $1$ for Allow None; $2$ for Allow A Few; $3$ for Allow Some; $4$ for Allow Many. For a more straightforward interpretation, we redefine the outcome variable to be binary: we let the outcome of the survey respondent be $1$ if she answers $3$ or $4$, indicating “support”; otherwise, we let her outcome be 0, indicating “oppose.” The treatment is also a binary variable, which equals 11 $(0)$ if the survey is about high-skilled (low-skilled) immigrants. We use a simple difference-in-means estimator to estimate the treatment effect of each country. The table below shows the point estimates and their standard errors. The point estimates speak to the difference between the percentage of people who support high-skilled immigrants and the percentage of people who support low-skilled immigrants. For example, 25.8% more survey respondents in Austria are more supportive of high-skilled immigrants, compared to low-skilled immigrants.
@@
Notes: Difference-in-difference estimates of the policy’s treatment effect by country. See text for details on how the outcome variable was constructed.
@@
$\max\left\{\frac{|\widehat{\tau}(X_{i})-\widehat{\tau}(X_{j})|}{\|X_{i}-X_{j}\|}\right\},\quad\forall i,j\in\mathcal{S}_{E}\cup\mathcal{S}_{P},$ (24)
@@
Notes: The dots in Panel 10(a) represent all possible pairs among the $15$ countries. The slope of each point connecting to the origin represents the value of the Lipschitz constant needed to explain the estimated treatment effects and site-level covariates observed for that pair of countries. The red dash line is the smallest $C$ that is needed to explain all the data, computed using equation 24. The slope of the blue dashed line represents $C(\mathscr{S})$, computed using 25. Panel 10(b) presents a histogram of $C(\mathscr{S})$ for all possible choices of experimental sites $\mathscr{S}$. The red dashed line is the smallest Lipschitz constant needed to explain the data, corresponding to the same red dashed line in panel 10(a).
@@ B.4 A Simple Example that Motivates Assumption 4
where $X_{i,s}\in\mathbb{R}$ is the observed unit-level covariate for individual $i$ in site $s$, $Z_{i,s}\in\mathbb{R}$ is the unobserved unit-level covariate, $\varepsilon_{i,s}\sim\mathcal{N}(0,\sigma_{\varepsilon,s}^{2})$ is an error term with $\sigma_{\varepsilon,s}^{2}>0$ and is independent of $\left(X_{i,s},Z_{i,s}\right)$, and $\beta,\gamma\in\mathbb{R}$ are the same across different sites. For simplicity, suppose that $X_{i,s}$ and $Z_{i,s}$ are jointly normal: $\left(X_{i,s},Z_{i,s}\right)^{\top}\sim\mathcal{N}(\mu_{s},\varSigma_{s})$, where $\mu_{s}\in\mathbb{R}^{2}$ and $\varSigma_{s}$ is positive definite.
Let $\bar{Y}_{s}:=\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}Y_{i,s}$ be the sample average of the observed outcome at site $s$ and $\bar{X}_{s}:=\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}X_{i,s},\bar{Z}_{s}:=\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}Z_{i,s}$ the observed and unobserved site-level aggregate covariates, respectively. Under the above assumptions, we have
@@
where $\mathbb{E}[\bar{Z}_{s}|\bar{X}_{s}]$ is the expectation of $\bar{Z}_{s}$ conditional on observed site-level aggregate covariate $\bar{X}_{s}$ and where $\sigma_{s}^{2}>0$. Then, Assumption 4 models a case for which the policy effect of interest is
@@
Furthermore, (26) implies that, conditional on $\bar{X}_{s}$, we have an unbiased and normal estimator for $\tau_{s}$. We may also calculate that $\mathbb{E}[\bar{Z}_{s}|\bar{X}_{s}]=\alpha_{s}$ for some $\alpha_{s}$ that depends on each site $s$. Then, for $^{\prime}\in\mathcal{S}, sss\neq s^{\prime}$, we have: